OpenStudy (anonymous):

Find the square roots of -3+4i

OpenStudy (earthcitizen):

square roots ?

OpenStudy (anonymous):

5.73205081 i according to google :/

OpenStudy (earthcitizen):

i do know if you convert to polar form, r = 5(radius),theta =126.87°(angle)

OpenStudy (anonymous):

I haven't learned polar form yet... It's using the (a^2-b^2)+2abi method.

OpenStudy (anonymous):

How do you factor a^4+3a^2+4=0?

OpenStudy (callisto):

i think you cannot factor it ?! (not include using the quadratic formula)

OpenStudy (earthcitizen):

with that you get four roots then

OpenStudy (anonymous):

When it's a^4-3a^2-4=0, it factors as (a^2-4)(a^2+1)..?

OpenStudy (earthcitizen):

@order (a^2-a+2) (a^2+a+2) = 0, is a = -3+j4 ?

OpenStudy (anonymous):

Not sure... does this have to do with my question?

OpenStudy (earthcitizen):

yep, yep

OpenStudy (anonymous):

(1 + 2i)^2 = -3+4i

OpenStudy (anonymous):

 solve that system to get a and b...